OpenStudy (mayaal):

giving a medals!please help! which of the folowing must be true for all values of x? I.(X+1)^2 is greater than or equal to x^2 II.(x-2)^2 is greater than or equal to 0 III.x^2 +1 is greater than or equal to 2x

OpenStudy (mayaal):

@jim_thompson5910 and @CrazyMad plz help

jimthompson5910 (jim_thompson5910):

are you able to graph each expression?

OpenStudy (mayaal):

this is a question on the psat so i don't think i will be able to use a graphing calculator.

jimthompson5910 (jim_thompson5910):

what does (x+1)^2 expand to?

OpenStudy (mayaal):

x^2 +2x +1,right?

OpenStudy (anonymous):

it is the 2nd option as anything in side the ( ) will be turned positive once it is squared

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

so \[\Large (x+1)^2 \ge x^2\] turns into \[\Large x^2+2x+1 \ge x^2\]

OpenStudy (mayaal):

mhmm.. i'm starting to get it

jimthompson5910 (jim_thompson5910):

we can then subtract x^2 from both sides to get \[\Large 2x+1 \ge 0\]

jimthompson5910 (jim_thompson5910):

what do you get once you solve for x?

OpenStudy (mayaal):

x> or equal to -1/2

jimthompson5910 (jim_thompson5910):

so \[\Large (x+1)^2 \ge x^2\] is only true if \[\Large x \ge -\frac{1}{2}\]

jimthompson5910 (jim_thompson5910):

therefore, inequality I is not true for all real numbers

OpenStudy (mayaal):

yes...

OpenStudy (mayaal):

ok...so do we do the same thing for all 3?

OpenStudy (aum):

ii) Whenever you square a real number or a real function it is always greater than or equal to zero: { f(x) }^2 >= 0 Therefore \((x-2)^2 \ge 0\) (always).

jimthompson5910 (jim_thompson5910):

agreed with aum on #2 squaring any real number will never result in a negative number

OpenStudy (aum):

iii)\[ (x-1)^2 \ge 0 \\ x^2 - 2x + 1 \ge 0 \\ x^2 + 1 \ge 2x \]

OpenStudy (mayaal):

so how do i know which one is true for all real numbers?

OpenStudy (aum):

ii) and iii)

OpenStudy (mayaal):

and can u explain why 3 isn't correct,again?

jimthompson5910 (jim_thompson5910):

\[\Large x^2 +1 \ge 2x\] \[\Large x^2 + 1 - 2x\ge 0\] \[\Large x^2 -2x + 1\ge 0\] \[\Large (x-1)^2 \ge 0\]

jimthompson5910 (jim_thompson5910):

(x-1)^2 is never negative, so it is either 0 or greater than 0

OpenStudy (mayaal):

x2+1≥2x x2+1−2x≥0 x2−2x+1≥0 (x−1)2≥0

OpenStudy (mayaal):

oops...sorry,lol.

jimthompson5910 (jim_thompson5910):

yeah it looks like you got it

OpenStudy (mayaal):

haha:) but, (x-1)^2 isn't even one of the options.

jimthompson5910 (jim_thompson5910):

it looks like (x−1)2≥0 is trying to say \[\Large (x-1)^2 \ge 0\]

jimthompson5910 (jim_thompson5910):

refresh the page if you get any crazy looking symbols

OpenStudy (mayaal):

the thing is,(x-1)^2 isn't one of the 3 options i gave u,so where did it come from?

jimthompson5910 (jim_thompson5910):

what are the options again?

OpenStudy (mayaal):

I.(X+1)^2 is greater than or equal to x^2 II.(x-2)^2 is greater than or equal to 0 III.x^2 +1 is greater than or equal to 2x

jimthompson5910 (jim_thompson5910):

I got \[\Large (x-1)^2 \ge 0\] starting from \[\Large x^2 +1 \ge 2x\]

jimthompson5910 (jim_thompson5910):

I show the steps above

OpenStudy (aum):

We have to prove or disprove x^2 +1 is greater than or equal to 2x x^2 + 1 >= 2x x^2 + 1 - 2x >= 0 (x-1)^2 >= 0 Any function square will always be greater than equal to zero for ALL real values of x. Therefore, (x-1)^2 >= 0 is always true which makes x^2 + 1 >= 2x because we derived (x-1)^2 >=0 from x^2 + 1 >= 2x

jimthompson5910 (jim_thompson5910):

you get (x-1)^2 from factoring x^2-2x+1

OpenStudy (mayaal):

ok guyz...so 3 is always true,right?

OpenStudy (aum):

Yes, 2 and 3 are always true.

OpenStudy (mayaal):

but 1 is false because it has to be >= 0,right?

jimthompson5910 (jim_thompson5910):

it's only true when x >= -1/2

jimthompson5910 (jim_thompson5910):

otherwise, it's false

jimthompson5910 (jim_thompson5910):

so overall, it's false for all real numbers

OpenStudy (mayaal):

so the final answer is:both 2 and 3 are true.

OpenStudy (mayaal):

@jim_thompson5910 ??

jimthompson5910 (jim_thompson5910):

yes only 2 and 3 are true

OpenStudy (mayaal):

thanku sooo much guyz.. what level of math r u in?